patrick is 3 years older than amy. in 5 years the sum if thier ages will be 31. how old are they now?
Answer by ptaylor(2198)

Let x=Amy's age
x+5=Amy's age in 5 years
x+3=Patrick's age
x+3+5=x+8=Patrick's age in five years
Now we are told the following:
x+5+x+8=31
2x+13=31 subtract 13 from each side
2x+13-13=31-13 collect like terms
2x=18 divide each side by 2
x=9------------------Amy's age
x+3=9+3=12---------------------Patrick's age
CK
(9+5)+(12+5)=31
14+17=31
31=31

Hope this helps---ptaylor